OpenStudy (anonymous):

6^11n+6=1,296? show work

OpenStudy (ranga):

Is n an exponent? 6^(11n) or 6^11 * n ?

OpenStudy (ranga):

or 6^(11n + 6) which seems more likely.

OpenStudy (ranga):

\[\Large 6^{11n + 6} = 1296\]

OpenStudy (ranga):

Is the last one the correct representation?

OpenStudy (anonymous):

yes @ranga

OpenStudy (ranga):

Okay. 1296 happens to be 6^4. So \[\Large 6^{11n + 6} = 1296 = 6^4\] Since the base is the same on the left hand side and right hand side, the exponents must be the same. So 11n + 6 = 4 solve for n.

OpenStudy (anonymous):

so whats the answer @ranga

OpenStudy (ranga):

I am not allowed to give out answers. 11n + 6 = 4 To solve for n, isolate n. Subtract 6 from both sides. Divide both sides by 11 to find n.

OpenStudy (anonymous):

11=-2?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

11n = -2 divide both sides by 11

OpenStudy (anonymous):

n=-2/11

OpenStudy (ranga):

Yes.